solve
8/z=9/z-1/z

Found 2 solutions by Alan3354, JBarnum:
Answer by Alan3354(69443)

8/z=9/z-1/z
----------
8/z = 8/z
z <> 0, all other values are valid

Answer by JBarnum(2146)

look at it does it matter what number z=? whether 1, 10 ,1890, -5,-78943, so it has to be Z=-infinity to +infinity